OpenStudy (spittman):

v=pir^2h solve for r

OpenStudy (3mar):

Is that is your form? \[v=\pi . r^2.h\] I think this is the volume of a right circular cylinder.

OpenStudy (3mar):

to solve for r: \[r^2=\frac{ v }{ \pi. h }\]

OpenStudy (3mar):

Take the square root: \[r=\sqrt{\frac{ v }{ \pi . h }}\]

OpenStudy (3mar):

just put the value of the volume and the height of the cylinder, you get directly the radius of the base (or the top) circle.

OpenStudy (mrnood):

@3mar The poster proably has not learned any methods that would help him with a similar question, from you writing the answer. It is better to guide him to ht emethod and let him DO the working himself with your help.

OpenStudy (3mar):

Thanks, MrNood. This requires the asker to be active, not out-of-response.